OpenStudy (anonymous):

3/x^2-x-20 = 2/x-5 + 1/x+4

OpenStudy (anonymous):

3/x^2-x-20 = 2/x-5 + 1/x+4 <===> 3/x^2-x-20 = (2x+8+x-5)/x^2-x-20 ===> 3= 3x+3 ===> x= 0

OpenStudy (anonymous):

3/(x²-x-20) = 2/(x-5) + 1/(x+4) 3/(x-5)(x+4) = 2/(x-5) + 1/(x+4) 3/(x-5)(x+4) = [2(x-4) + 1(x-5)]/(x-5)(x+4) 3 = 2(x-4) + 1(x-5) 3 = 2x - 28 + x - 5 3 = 3x - 33 1 = x - 11 x = 12